OpenStudy (zenmo):

Linear and Nonlinear systems of equations. Solve the system with substitution. -(1/2)x+y=-(5/2) x^2+y^2=25

OpenStudy (astrophysics):

What's the question? Solve for x and y?

OpenStudy (astrophysics):

Oh I see it says substitution :P

OpenStudy (zenmo):

My bad :X, updated the question.

OpenStudy (astrophysics):

Alright, so lets solve for y in the linear equation and plug it into the equation of circle they gave you

OpenStudy (astrophysics):

Any idea how?

OpenStudy (astrophysics):

\[y = -\left( \frac{ 5 }{ 2 } \right) + \left( \frac{ 1 }{ 2 } \right)x\] right?

OpenStudy (zenmo):

Yea sorry, I'm still here. Ok

OpenStudy (astrophysics):

Now lets plug that into the second equation, can you do that?

OpenStudy (zenmo):

Ok, I gotten \[x ^{2}+(\frac{ 1 }{ 2 }x)-\frac{ 55 }{ 2 } =0\]. Then we use the quadratic equation right to find the two X's?

OpenStudy (astrophysics):

\[x^2+ \left( -\frac{ 5 }{ 2 }+\frac{ 1 }{ 2 }x \right)^2=25\] this is what you get when you plug it in right

OpenStudy (astrophysics):

Not sure what you're doing

OpenStudy (astrophysics):

\[\left( -\frac{ 5 }{ 2 } +\frac{ 1 }{ 2 }x\right)\left( -\frac{ 5 }{ 2 } +\frac{ 1 }{ 2 }x\right)\] you need to expand this :), and no worries

OpenStudy (zenmo):

\[x ^{2}-\frac{ 25 }{ 4 }+\frac{ 1 }{ 4 }x ^{2}\]. That next step?

OpenStudy (astrophysics):

You should get \[\frac{ x^2 }{ 4 }-\frac{ 5x }{ 2 }+\frac{ 25 }{ 4 }\]

OpenStudy (astrophysics):

\[x^2+\frac{ x^2 }{ 4 }-\frac{ 5x }{ 2 }+\frac{ 25 }{ 4 } = 25\] this is what you should have now

OpenStudy (zenmo):

\[x ^{2}-\frac{ 25 }{ 4 }+\frac{ 1 }{ 4 }x ^{2} -> (\frac{ 4 }{ 4 }x ^{2}+\frac{ 1 }{ 4 }x ^{2})=5x ^{2}-\frac{ 25 }{ 4 }=25\]

OpenStudy (astrophysics):

Check your work

OpenStudy (astrophysics):

And signs

OpenStudy (zenmo):

How do you get \[\frac{ x ^{2} }{ 4 }-\frac{ 5x }{ 2 }+\frac{ 25 }{ 4 }\]?

OpenStudy (astrophysics):

\[\left( -\frac{5 }{ 2 }+\left( \frac{ 1 }{ 2 }x \right) \right)^2 \implies (\left( -\frac{5 }{ 2 }+\left( \frac{ 1 }{ 2 }x \right) \right)\left( -\frac{5 }{ 2 }+\left( \frac{ 1 }{ 2 }x \right) \right)\]

OpenStudy (astrophysics):

Yeah it's a messy question :P don't worry all I did was "foil"

OpenStudy (zenmo):

Yea I'm working on it still, the answer on the book shows (-3,-4) and (5,0). So, I'm figuring out how to get those at the moment.

OpenStudy (astrophysics):

We will have to factor we're almost at that process

OpenStudy (astrophysics):

But tag me if you have anymore troubles good luck!

OpenStudy (zenmo):

Yea, thanks for your time.

OpenStudy (astrophysics):

Np!